OpenStudy (mwhit2101):

Canadian Snow Geese migrate more than 4,800 kilometers from their winter home in the Southwest United States to the Canadian arctic. One flock was known to migrate 600 kilometers in 7.5 hours. Write a direct variation equation for the distance flown at any time. How long will it take the geese to travel 4,800 kilometers at this rate?

OpenStudy (bibby):

set up the proportion keeping the units aligned.\[\huge \frac{ 600km }{ 7.5hr } = \frac{ 4800km }{ xhr }\]

OpenStudy (bibby):

to find the variation you can either cross multiply or just find the relation between 4800 and 600.

OpenStudy (mwhit2101):

How?

OpenStudy (mwhit2101):

You mean cross multiply like 4800*7.5?

OpenStudy (bibby):

How do you get from 600 to 4800? 600x = 4800

OpenStudy (bibby):

that works too

OpenStudy (mwhit2101):

so 36000...

OpenStudy (mwhit2101):

4800*7.5=36000 i believe

OpenStudy (bibby):

you didn't multiply across the other diagonal

OpenStudy (mwhit2101):

to get from 600 to 4800 you would divide it which gives you 8

OpenStudy (mwhit2101):

Well i dont know the hours so i cant...

OpenStudy (mwhit2101):

unless i just put 600x idk

OpenStudy (bibby):

that's the whole point of algebra... You use x for an unknown variable you're solving for. 600x=36000

OpenStudy (mwhit2101):

ohh

OpenStudy (mwhit2101):

ok so

OpenStudy (bibby):